OpenStudy (anonymous):

The size of bacteria triples every 12 years. If the piece of bacteria was 25mm in 1990, within what year will the bacteria be 129 mm in size. Options are: 2000, 2002, 2008, and 2014 show work please

OpenStudy (anonymous):

define: bacteria grows according to the function g(k) = a*3^(k/12) where a is initial length and k is number of years since bacteria was at initial length. we solve for 129 = 25*3^(k/12) ; k = 17.9 approx = 18 years after it was 25 mm long. Thus the answer is 1990 +18 = 2008